INAF queue observing, 2016 Jan 11-12

Observer: R. Carini, A. Marchetti
Telescope Operator: Steve Allanson
Support Astronomer: B. Rothberg on call
Instrument:

Summary

At the sunset the sky is clear, humidity 20%.

There are some technical problems (as rotator and elevation axes..) in opening the dome

02:00 the problems are resolved, the TO take some test.

Night duration (12\xB0 el twilight) 19:00-6:00 ( 02:00am-01:00pm UT) = 11 h 00m

Observing time: 10 hrs 30 m

Weather time loss: 0 hrs * *(0%)

technical time loss 30 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
Saracco M004XLSms2 red grating 1h seeing var variable 1.0''-1.5'', with peaks up to 1.6'' or 1.8''
Antoniucci v1118ori dual grating +ima 30 m completed
Antoniucci v1647ori dual grating+ ima 30m to be repeated
Marconi IZw18 V, I 10 m ob1 done
Bellazzini C 086.4+31.8 g,r 0.2 h completed
Annibali DDO68 g, r 39 m the second ob it has to be completed (5/8 are good)
Bellazzini I 143.7+12.9 g,r 0.2 h completed
Bellazzini J 147.0+07.1 g,r 0.2 h completed
Bellazzini K 183.0+04.4 g,r 0.2 h completed
Bellazzini L 195.9+06.9 g,r 0.2 h completed
Bellazzini M 196.6+06.5 g,r 0.2 h completed
Std. Phot. SA104 blue:SDT_U,B,V,g; red: I, r, V,F972
Sky flat V bessel in red channel and B Bessel in the blue channel    
Bias LBC      

Details

01:23 Wake up MODS
  • take test images mods1r.20160108.0101,mods1b.20160108.0081

02:17 begin the saracco program
  • acqMODS mods.4.XLSms2.acq mods1r.20160112.0001-0004

02:41 execMODS mods.4.XLSms2_3.obs mods1r.20160112.0005-0008
  • seeing very variable 1.1''-1.6''
  • 02:55 better seeing, variable between 1.0'' and 1.3''
  • 03:08 seeing variable 1.2''-1.5''
  • 03:30 mean seeing 1.5''
  • 03:38 seeing 1.2''-1.4''

03:54 the seeing is too variable and it is more than 1.2'', so out of spec, we slew to Antoniucci target V1118ori( back up program)
  • acqMODS v1118ori_ls.acq mods1r.20160112.0009-0011
  • execMODS v1118ori_ls.obs mods1r.20160112.0012-0015 mods1b.20160112.0001-0004
  • 04:07 seeing 1.4''-1.6'' mods1r.20160112.0016-001

04:37 target V1647 ori
  • acqMODS v1647ori_ls.acq mods1r.20160112.0016-0019
  • execMODS v1647ori_ls.obs mods1r.20160112.0020-0023 mods1b.20160112.0005-0008 the blue images are very very faint, wrong posangle , to be repeat
  • mean seeing 1.5''
  • execMODS v1647ori.img mods1r.20160112.0024-0026, mods1b.0009-0011 the images present the effect of reflection

05:46 come back to V1118ori to run imaging
  • execMODS v1118ori.img mods1r.20160112.0027-0029, mods1b.0012-0014, the target is saturated, we take one image with texp=25 sec mods1r.0030, mods1b.0015, the blue is saturated , the red almost good, we take another image with texp=20 sec mods1r.0031, mods1b.0016 are saturated again, try texp=10sec, mods1r.0032, mods1b.0017, the blue has some pixel saturate, the red is ok, we reduce the texp for the blue t=5 sec. now it's ok mods1r.0033, mods1b.0018 we take other two image in the blue and one in the red, as asked by the PI. mods1r.0034, mods1b.0019-0020
  • seeing 1.0'' -1.3''

06:14 slew to the stellar standard GD71 (red + dual grating) * acqMODS gd71.acq mods1r.0035-0036 * execMODS gd71.obs mods1r.0037-00, mods1b.0021 * seeing about 0.9-1.0''

06:57 the seeing now is better (1''-1.2'') we pass to LBC

07:15 begin Marconi program
  • /dofpia, /x2, /first
  • copointing IZw18_copointing.ob
  • 07:30 problem with lbcrangebal with device, close and start again ds9 and iraf
  • 07:40 copointing again with success
  • 07:43 start the science IZw18_1.ob
  • seeing 1.0''

08:02 slew to Bellazzini program, target C
  • dofpia, /x2
  • copointing 2015B_1_C_pointing.ob, problem with the copointing, ds9 do not speck with iraf,close and start again ds9 and iraf
  • copointing again
  • 08:20 start the science 2015B_1_C.ob
  • seeing 1''-1.2''

08:40 the seeing is out of spec to continue Bellazzini program, we slew to Annibali
  • dofpia, /x
  • copointing
  • 08:54 we send the calibration ob given us from the PI
  • 08:58 send scientific ob DDO68_1_new.ob
  • seeing 0.8''- 1.0''
  • the last 4 images have stars a little bit elongated, we run again the dofpia and the copointing, the images are not better, run the dofpia again with double time exposure

09:54 send the second scientific ob DDO68_2_new.ob
  • seeing 0.8''
  • the first image is good (lbcr.20160112.095436, lbcb.20160112.095441)
  • from the the image 2 the sources are elongated, the target is at zenith and the guider has problems. at a high elevation (less than 1) the LBCs are rotating really fast its hard to maintain good guiding the images are elongated in both, we have to change target

10:15 slew to target I of Bellazzini
  • dofpia, copointing 2015B_1_I_copoint.ob
  • run science 2015B_1_I.ob
  • seeing 0.9''
  • 10:28 during the acquisition of the first image, the tracker gives error
  • seeing 1.0''-1.1''

10:42 slew to J target
  • the TO restarts the tracker
  • 10:48 dofpia, copointing 2015B_1_J_copoint.ob
  • 11:00 run the scientific ob 2015B_1_J.ob
  • seeing 0.8''-1.0''

11:15 slew to K target
  • technical error
  • 11:24 the TO resolves the problem, run dofpia again and copointing 2015B_1_K_pointing.ob
  • 11:37 run science 2015B_1_K.ob
  • seeing about 1.0''

11:51 slew to L target
  • dofpia, copointing 2015B_1_L_pointing.ob
  • 12:02 run science 2015B_1_L.ob
  • seeing about 0.8''

12:17 slew to M target
  • dofpia, this target is very near to L, so it is useless to run the copointing
  • 12:22 run science 2015B_1_M_pointing.ob
  • seeing 0.8''-0.9'''

12:38 slew to Annibali target
  • dofpia and copointing DDO68_copointing.ob
  • 12:48 run the scientific ob DDO68_2_new.ob
  • seeing 0.6''-0.7''
  • the first image is ok in both channel (lbcr. ,lbcb.20160112.124457), the second imaging is smearing, the TO restart LBC.
  • the elongation is due to a tracker error, the images were not actually guiding.
13:00 run DDO68_2_new.ob
  • seeing 0.8

13:10 stop the science, airmass out of spec. the first three images are good, the other have elongated sources.

13:20 run the photometric standard

13:30 skyflat

14:15 bias LBC

-- %USERSIG{RobertaCarini - 2016-01-12}%

Comments

 
Topic revision: r13 - 13 Jan 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback